OpenStudy (anonymous):

Which is the best estimate of x based on rounding the constants and coefficients in the equation to the nearest integer? 3.1x + 5.2 = 9.9

OpenStudy (anonymous):

You'd round 3.1 down to 3, 5.2 to 5, and 9.9 to 10. 3x+5=10 -5 on both sides 3x=5 divide by 5 x= 5/3. I don't know what your options are, but that's the solution.